Jack and Jill ran up a hill to fetch a pail of water. They ran 100m to get to the top of a 30m tall hill. What is the degree of incline of the hill?
1
Expert's answer
2020-05-21T17:59:38-0400
"\sin\theta=\frac{P}{H}=\frac{30}{100}=0.3\\or\\\theta = \sin^{-1}(0.3)=17.45\degree"

Inclination = "17.45\degree"